You just read a report that contained the following boundaries for a 90% confidence interval bassed on a sample proportion of .36: 0.2284 and 0.4916. What was the associated sample size?
Answer #1

Margin of error = (0.4916 - 0.2284)/2 = 0.1316

For 90% confidence, z = 1.645

p = 0.36 So q = 1 - 0.36 = 0.64

Hence,

On solving above for n, we will get:

n = 36